### v2

# Absence of twisting for non-trivial discrete torsion

Abstract.
We study discrete torsion for the nn–torus with finite symmetry group $G$ from the Dijkgraaf–Witten viewpoint. A class in $H^{n}(G,U(1))$ assigns a phase to each flat GG–bundle, equivalently to each commuting nn–tuple in $G$ up to conjugation. We introduce the subgroup $\operatorname{Utw}^{n}(G)\subseteq H^{n}(G,U(1))$ of untwisted classes, those whose Dijkgraaf–Witten phases are trivial on all commuting tuples, and derive a universal coefficient exact sequence involving this invariant. In degree $2$ this recovers the Bogomolov multiplier / unramified Brauer group. We implement algorithms for computing $\operatorname{Utw}^{n}(G)$ and corresponding torus partition functions, and report on computations for families of finite subgroups of $\operatorname{SU}(4)$.
The author was partially supported by ARIS, Grant numbers J1-50001, P1-0222.
1. Introduction
Discrete torsion is one of the fundamental tools used for describing string orbifolds. It was first discovered by Vafa [28] as a somewhat mysterious way of choices of phases corresponding to twisted vectors of string orbifold partition functions. These describe the effects of the orbifold group action on the $B$-field in string theory [1], [26]. Discrete torsion also generalizes orbifold projections of D-brane probes at space-time singularities. In a series of papers, Sharpe clarified the mathematical background of discrete torsion, see [26, 27] for overview.
In a two-dimensional orbifold CFT with finite orbifold group $G$, discrete torsion is specified by a cohomology class $\omega]\in H^{2}(G,U(1))$. The latter group is also known as the Schur multiplier of $G$. A representative 2-cocycle $\omega\in Z^{2}(G,U(1))$ assigns a phase $W_{\omega}(g,h)$ to each commuting pair $(g,h)$ of group elements, which enters the one-loop (torus) orbifold partition function as
(1.0.1) $Z^{\text{orb}}_{\omega}(T^{2})\;=\;\frac{1}{|G|}\sum_{\begin{subarray}{c}g,h\in G\\ gh=hg\end{subarray}}W_{\omega}(g,h)\,Z_{g,h}(T^{2}),$
where $Z_{g,h}(T^{2})$ is the contribution from the (g,h)(g,h)–twisted sector. For the standard discrete–torsion construction one takes
(1.0.2) $W_{\omega}(g,h)\;=\;\frac{\omega(g,h)}{\omega(h,g)}\,.$
Beyond its original CFT definition, discrete torsion admits a clean modern formulation. It is a choice of topological coupling for a finite symmetry group $G$, encoded by a class in $H^{n}(G,U(1))$, and realized by coupling the theory to the corresponding Dijkgraaf–Witten (DW) topological gauge theory [7]. On a closed nn–manifold $M$, this coupling contributes a phase obtained by evaluating $f_{A}^{*}[\omega]$ on $[M]$, where $f_{A}:M\to BG$ classifies the flat GG–bundle AA.
A subtlety, which is easy to miss if one focuses only on the abstract group $H^{n}(G,U(1))$, is that different cohomology classes can become physically indistinguishable on a given class of backgrounds. For the nn–torus $T^{n}$, the DW weight $W_{\omega}(\mathbf{g})$ depends only on the evaluation of $\omega]$ on the fundamental class through the classifying map associated to a commuting nn–tuple $\mathbf{g}$. This leads to a natural notion of untwisted discrete torsion. These are the cohomology classes whose associated phases are trivial on all commuting tuples. Such classes are invisible to the torus partition function, and more generally to any observable that factors through the DW holonomy on flat GG–bundles on $T^{n}$.
The central theme of this paper is that “invisibility on tori” has a robust group-theoretic meaning, and can be computed efficiently. In degree $2$ it recovers the unramified Brauer group, also known as the Bogomolov multiplier [2], and admits a dual description which makes large-scale computation feasible. In higher degrees we introduce and study the groups $\operatorname{Utw}^{n}(G)$ of untwisted nn–dimensional discrete torsion, prove a universal coefficient sequence tailored to this setting, and explain how the obstruction to twisting can be read off from an explicit quotient $H_{0n}(G,\mathbb{Z})$ of $H_{n}(G,\mathbb{Z})$. This provides a homological and computational bridge between discrete torsion phases and the topology of the moduli space of flat GG–connections on $T^{n}$. We provide an implementation [23] of the constructions in publicly available package HAP [10] of the computer algebra system GAP [9].
On the physics side, discrete torsion is often invoked as a genuine deformation parameter of an orbifold, yet in concrete families the “available” torsion in $H^{n}(G,U(1))$ may fail to change torus amplitudes at all. Understanding when this happens clarifies which orbifolds admit genuinely distinct topological couplings and which ones exhibit a rigidity phenomenon. Our results supply a systematic diagnostic for when discrete torsion is physically effective. This is precisely the complement of $\operatorname{Utw}^{n}(G)$, and it controls whether twisting is genuinely present in torus amplitudes.
2. Partition function of the $n$-torus
Let $\mathcal{T}$ be a $d$-dimensional QFT (or CFT) with finite global symmetry group $G$, and let $\omega]\in H^{d}(G,U(1))$ encode a topological $G$-action (discrete torsion). For a closed oriented $d$-manifold $M$, a flat $G$-bundle AA on $M$ is equivalently a homomorphism $f_{A}:\pi_{1}(M)\to G$, or a map $f_{A}:M\to BG$ up to homotopy. The Dijkgraaf–Witten topological gauge theory [7] with action $\omega$ has partition function
(2.0.1) $Z_{\omega}^{\mathrm{DW}}(M)\;=\;\frac{1}{|G|}\sum_{[A]\ \mathrm{flat}}\exp\!\bigl(2\pi i\,S_{\omega}(A)\bigr)\;=\;\frac{1}{|G|}\sum_{[A]\ \mathrm{flat}}\bigl\langle f_{A}^{*}[\omega],[M]\bigr\rangle,$
where $f_{A}^{*}:H^{d}(G,U(1))\to H^{d}(\pi_{1}(M),U(1))$ is the map induced by $f_{A}$, and $S_{\omega}(A)$ is the evaluation of $f_{A}^{*}[\omega]\in H^{d}(M,U(1))$ on the fundamental class $[M]$.
The orbifold of $\mathcal{T}$ by $G$ with discrete torsion $\omega]$ is obtained by coupling $\mathcal{T}$ to this topological DW gauge theory and summing over flat $G$-bundles. Its partition function on $M$ is
(2.0.2) $Z^{\mathrm{orb}}_{\omega}(M)\;=\;\frac{1}{|G|}\sum_{[A]\ \mathrm{flat}}\bigl\langle f_{A}^{*}[\omega],[M]\bigr\rangle\,Z_{\mathcal{T}}(M;A),$
where $Z_{\mathcal{T}}(M;A)$ denotes the partition function of the original theory $\mathcal{T}$ on $M$ in the background of the flat $G$-connection AA.
We now specialize to $M=T^{n}=S^{1}\times\cdots\times S^{1}$. A flat GG–bundle on $T^{n}$ is determined by its holonomies along the $n$ circle factors, i.e. by a commuting nn–tuple
(2.0.3) $\mathbf{g}=(g_{1},\dots,g_{n})\in G^{n},\qquad g_{i}g_{j}=g_{j}g_{i}.$
Denote
$X_{n}(G)=\{(g_{1},\dots,g_{n})\in G^{n}\mid g_{i}g_{j}=g_{j}g_{i}\text{ for all }i,j\}.$
For a commuting tuple $\mathbf{g}=(g_{1},\dots,g_{n})$, let $f_{\mathbf{g}}:T^{n}\to BG$ be the corresponding classifying map. The DW weight of this background is
(2.0.4) $W_{\omega}(\mathbf{g})\;:=\;\big\langle f_{\mathbf{g}}^{*}[\omega],[T^{n}]\big\rangle\;\in U(1),$
so that
(2.0.5) $Z^{\mathrm{DW}}_{\omega}(T^{n})\;=\;\frac{1}{|G|}\sum_{\mathbf{g}\in X_{n}(G)}W_{\omega}(\mathbf{g})$
and
(2.0.6) $Z^{\mathrm{orb}}_{\omega}(T^{n})\;=\;\frac{1}{|G|}\sum_{\mathbf{g}\in X_{n}(G)}W_{\omega}(\mathbf{g})Z_{\mathcal{T}}(T^{n};\mathbf{g}).$
To obtain an explicit formula for $W_{\omega}(\mathbf{g})$, represent $T^{n}$ as the unit cube $0,1]^{n}$ with opposite faces identified, and equip it with the standard orientation $\mathrm{d}x_{1}\wedge\cdots\wedge\mathrm{d}x_{n}$. There is a canonical decomposition of the cube into $n!$ nn–simplices indexed by permutations $\sigma\in S_{n}$:
(2.0.7) $\Delta_{\sigma}\;=\;\big\{0\leq t_{1}\leq\cdots\leq t_{n}\leq 1\big\}\ni x\;=\;t_{1}e_{\sigma(1)}+\cdots+t_{n}e_{\sigma(n)},$
where $e_{i}$ are the standard basis vectors of $\mathbb{R}^{n}$. Each $\Delta_{\sigma}$ inherits an orientation; its oriented volume form is $\operatorname{sgn}(\sigma)\,\mathrm{d}x_{1}\wedge\cdots\wedge\mathrm{d}x_{n}$, so the relative orientation of $\Delta_{\sigma}$ with respect to $T^{n}$ is $\operatorname{sgn}(\sigma)$.
In the DW construction one chooses a gauge in which a distinguished vertex of the triangulation is labeled by the identity of $G$, and the edge along the ii–th coordinate direction is labeled by the group element $g_{i}$. For an oriented nn–simplex whose edges follow the ordered sequence of coordinate directions $(\sigma(1),\dots,\sigma(n))$, the local DW weight is precisely
(2.0.8) $\omega\big(g_{\sigma(1)},\dots,g_{\sigma(n)}\big)$
(or its inverse, depending on orientation). Therefore, the total DW weight for the commuting nn–tuple $(g_{1},\dots,g_{n})$ is given by the product over all nn–simplices,
(2.0.9) $W_{\omega}(g_{1},\dots,g_{n})\;=\;\prod_{\sigma\in S_{n}}\omega\big(g_{\sigma(1)},\dots,g_{\sigma(n)}\big)^{\operatorname{sgn}(\sigma)}.$
This can be compared with [13], and is the natural nn–dimensional generalization of the familiar two-dimensional discrete torsion phase $\omega(g,h)/\omega(h,g)$ and of the 3-torus expression
(2.0.10) $W_{\omega}(g,h,k)\;=\;\frac{\omega(g,h,k)\,\omega(h,k,g)\,\omega(k,g,h)}{\omega(g,k,h)\,\omega(h,g,k)\,\omega(k,h,g)},$
which appears as eq. (6.35) in [7].
Finally, one checks that (2.0.9) depends only on the cohomology class $\omega]\in H^{n}(G,U(1))$ and not on the choice of representative: if $\omega^{\prime}=\omega\,\delta\alpha$ with $\alpha$ an (n1)(n{-}1)–cochain, then the contribution of $\delta\alpha$ to $W_{\omega^{\prime}}(g_{1},\dots,g_{n})$ is the evaluation of a coboundary on the fundamental cycle $T^{n}]$, which is trivial. Equivalently, the extra factors from $\delta\alpha$ can be organized, using the cocycle identity, into pairs that cancel along shared (n1)(n{-}1)–faces of adjacent simplices.
We also observe that $G$ acts diagonally on the set of all commuting $n$-tuples in $G$. Conjugate tuples give the same bundle, compare this with [24]. So in practice one sums (2.0.5) and (2.0.6) over conjugacy class representatives of commuting $n$-tuples, with suitable weights, which can be computed by looking at stabilizer group sizes. We refer to Appendix A for further details.
Set
$\operatorname{Utw}^{n}(G)=\{[\omega]\in H^{n}(G,U(1))\mid W_{\omega}(\mathbf{g})=1\hbox{ for all }\mathbf{g}\in X_{n}(G)\}.$
The set $\operatorname{Utw}^{n}(G)$ is the set of all untwisted $n$-dimensional discrete torsion. We say that the group $G$ is $n$-untwisted if and only if $\operatorname{Utw}^{n}(G)=H^{n}(G,U(1))$. Of course, we are only interested in the case when $H^{n}(G,U(1))$ is non-trivial, so this will be assumed throughout the paper. Note that the untwisted groups are precisely the $2$-untwisted groups.
3. Untwisted discrete torsion and unramified Brauer groups
Let $\omega\in Z^{2}(G,U(1))$. An element $g\in G$ is said to be $\omega$-regular [17, p. 94] if
$\omega(g,h)=\omega(h,g)\hbox{ for all }h\in C_{G}(g).$
Again, one readily sees that the definition of $\omega$-regular elements only depends on the cohomology class of $\omega$. The definition clearly implies
$\operatorname{Utw}^{2}(G)=\{[\omega]\in H^{2}(G,U(1))\mid\hbox{ every element of }G\hbox{ is }\omega\hbox{-regular}\}.$
Davydov [6] denotes the right-hand side subgroup by $B(G)$. We observe that the group $\operatorname{Utw}^{2}(G)$ had been discovered independently multiple times. Bogomolov [2] showed that $\operatorname{Utw}^{2}(G)$ is isomorphic to the unramified Brauer group of $\mathbb{C}(V)^{G}/\mathbb{C}$, where $V$ is a representation of $G$. Based on that, $\operatorname{Utw}^{2}(G)$ is also known as the Bogomolov multiplier of $G$. This group has applications in Noether’s problem in invariant theory [25]. Independently, the group $\operatorname{Utw}^{2}(G)$ was also studied by Higgs [12], he denoted it by $H^{2}_{E}(G,U(1))$. A more general version was described in [14]. Very recently, Kobayashi and Watanabe [18] described the role of the elements of $\operatorname{Utw}^{2}(G)$ as labels of (1+1)D SPT phases that escape standard string-order diagnostics, and—after gauging—produce distinct gapped phases in which the non-invertible symmetry $\mathrm{Rep}(G)$ is completely broken.
We develop a method for computing the cocycles that represent the elements of $\operatorname{Utw}^{2}(G)$. We refer to Appendix A for further details. On the other hand, it is computationally much more suitable to use a dual description of the group $\operatorname{Utw}^{2}(G)$ for a given group $G$, see [22]. One first defines the non-abelian exterior square $G\wedge G$ of the group $G$ to be the group generated by the symbols $g\wedge h$, where $g,h\in G$, subject to the following relations:
$gg_{1}\wedge h$ $=(gg_{1}g^{-1}\wedge ghg^{-1})(g\wedge h)$
$g\wedge hh_{1}$ $=(g\wedge h)(hgh^{-1}\wedge hh_{1}h^{-1})$
$g\wedge g$ $=1$
for all $g,g_{1},h,h_{1}\in G$. The group $G\wedge G$ was first introduced by Miller [21], and later generalized by Brown and Loday [4] in connection with van Kampen theorem for diagrams of spaces. The commutator map $\kappa:G\wedge G\to G$ defined via $\kappa(g\wedge h)=ghg^{-1}h^{-1}$ is a homomorphism. Denote its kernel by $M(G)$. Miller [21] showed that $M(G)$ is isomorphic to the second integral homology group $H_{2}(G,\mathbb{Z})$. Let $M_{0}(G)=\langle x\wedge y\mid x,y\in G,\,xy=yx\rangle$. Then we have the following [22]:
Theorem 3.1.
Let $G$ be a group. Then $\operatorname{Utw}^{2}(G)$ is isomorphic to $\operatorname{Hom}(B_{0}(G),U(1))$, where $B_{0}(G)=M(G)/M_{0}(G)$. When $G$ is finite, we have a (non-natural) isomorphism $\operatorname{Utw}^{2}(G)\cong B_{0}(G)$.
Theorem 3.1 provides a description of the isomorphism type of $\operatorname{Utw}^{2}(G)$, which is suitable for efficient computations. An algorithm for computing $B_{0}(G)$ was first developed in [22], and subsequently improved in [14] for the case of finite solvable groups. An independent implementation is a part of the homological algebra package HAP [10] of the computer algebra system GAP [9].
It is clear that $G$ admits an untwisted discrete torsion if and only if $B_{0}(G)$ is non-trivial. It turns out [15] that these groups are, contrary to the previous folk-lore belief, quite abundant. Furthermore, $G$ is an untwisted group if and only if $B_{0}(G)=M(G)$. Groups $G$ with the latter property were also considered by Cameron and Kuzma [5] in connection with relationship between commuting and deep commuting graphs of groups.
It is known [16] that if $G_{1}$ and $G_{2}$ are finite groups, then $\operatorname{Utw}^{2}(G_{1}\times G_{2})\cong\operatorname{Utw}^{2}(G_{1})\times\operatorname{Utw}^{2}(G_{2})$. There is a well established fact [17, p. 46] that $H^{2}(G_{1}\times G_{2},U(1))\cong H^{2}(G_{1},U(1))\times H^{2}(G_{2},U(1))\times(G_{1}\otimes G_{2})$, where $G_{1}\otimes G_{2}$ is the tensor product of the largest abelian quotients of $G_{1}$ and $G_{2}$. We conclude the following:
Proposition 3.2.
Let $G_{1}$ and $G_{2}$ be untwisted groups of coprime orders. Then $G_{1}\times G_{2}$ is also untwisted.
We report on examples of untwisted groups in Section 5. As another example, we consider finite subgroups $G$ of $\operatorname{SU}(n)$, $n=2,3$. We refer to the notation of [20], and we note that the calculations of $H^{2}(G,U(1))$ done in [8]. Here we show that the subgroup $\operatorname{Utw}^{2}(G)$ is always trivial.
If $G$ is a finite subgroup of $\operatorname{SU}(2)$, then $H^{2}(G,U(1))=1$ [8, Corollary 1]. So we consider finite subgroups of $\operatorname{SU}(3)$. Here we apply the following facts [2, 22]:
Proposition 3.3.
Let $G$ be a finite group.
  1. (1)
    If $G$ is abelian, then $\operatorname{Utw}^{2}(G)$ is trivial.
  2. (2)
    If $S$ is a Sylow $p$-subgroup of $G$, then the $p$-part of $\operatorname{Utw}^{2}(G)$ embeds into $\operatorname{Utw}^{2}(S)$.
  3. (3)
    If $G$ has an abelian normal subgroup $A$ with $G/A$ cyclic, then $\operatorname{Utw}^{2}(G)$ is trivial.
  4. (4)
    If $G$ is a group of order $2^{n}$ with $\operatorname{Utw}^{2}(G)$ non-trivial, then $n\geq 6$.
  5. (5)
    If $G$ is a group of order $p^{n}$, $p>2$, with $\operatorname{Utw}^{2}(G)$ non-trivial, then $n\geq 5$.
In Table 1 we list the families of finite subgroups of $\operatorname{SU}(3)$ as in [20], with reasons implying the triviality of the Bogomolov multiplier. Note that if $G$ is a group of order $2^{n_{0}}3^{n_{1}}\cdots$ and if $n_{0}<6$ and $n_{i}<5$ for all $i\geq 1$, then the items 2, 4 and 5 of Proposition 3.3 give the result. This quickly rules out the last seven rows of the table. The remaining non-obvious cases are the families $C(n,a,b)$ and $D(n,a,b;d,r,s)$. The former can be decomposed as $C(n,a,b)\cong(\mathbb{Z}/m_{1}\times\mathbb{Z}/m_{2})\rtimes\mathbb{Z}/3$ for some $m_{1}$ and $m_{2}$ depending on $n$, $a$ and $b$, see [20]. Therefore, $C(n,a,b)$ has an abelian normal subgroup with cyclic quotient, hence $\operatorname{Utw}^{2}(C(n,a,b))$ is trivial by Proposition 3.3, 3. Similarly, we have $D(n,a,b;r,s)\cong(\mathbb{Z}/n_{1}\times\mathbb{Z}/n_{2})\rtimes S_{3}$ for some $n_{1}$ and $n_{2}$. From here we deduce that all Sylow $p$-subroups of $D(n,a,b;r,s)$ are either abelian, or of the form $A_{(2)}\rtimes\mathbb{Z}/2$ or $A_{(3)}\rtimes\mathbb{Z}/3$, where $A_{(q)}$ is an abelian $q$-group. This shows the result by Proposition 3.3, 1, 2 and 3.
Group Reason for triviality of $\operatorname{Utw}^{2}(G)$
$\mathbb{Z}/m\oplus\mathbb{Z}/n$ Proposition 3.3, 1
Finite subgroups of $SU(2)$ [8]
$C(n,a,b)$ Proposition 3.3, 3
$D(n,a,b;d,r,s)$ Proposition 3.3, 1, 2, 3
$\Sigma(60)\cong A_{5}$ Proposition 3.3, 2, 4, 5
$\Sigma(60)\times\mathbb{Z}/3$ Proposition 3.3, 2, 4, 5
$\Sigma(168)\cong\operatorname{PSL}(2,7)$ Proposition 3.3, 2, 4, 5
$\Sigma(36\times 3)$ Proposition 3.3, 2, 4, 5
$\Sigma(72\cdot 3)$ Proposition 3.3, 2, 4, 5
$\Sigma(216\cdot 3)$ Proposition 3.3, 2, 4, 5
$\Sigma(360\cdot 3)$ Proposition 3.3, 2, 4, 5
Table 1. Finite subgroups of $\operatorname{SU}(3)$ according to [20].
4. Higher-dimensional discrete torsion
Here we develop the machinery for computing the groups $\operatorname{Utw}^{n}(G)$. In order to do this, we first consider general cohomology of $G$ with coefficients in a trivial $G$-module AA. In general, we write the operation in $A$ additively, unless when $A=U(1)$; in that case we use the multiplicative notation. Denote by $F_{\bullet}\to\mathbb{Z}$ the normalized bar resolution of $\mathbb{Z}$ as a left $\mathbb{Z}G$–module, and set
$C_{\bullet}=C_{\bullet}(G):=\mathbb{Z}\otimes_{\mathbb{Z}G}F_{\bullet}\qquad(C_{0}=\mathbb{Z}).$
For $g_{1},\dots,g_{n}\in G$ we identify the element $1\otimes[g_{1}|\cdots|g_{n}]\in C_{n}$ with $g_{1}|\cdots|g_{n}]$ and suppress the tensor sign. We have the differential maps $\partial_{n}:C_{n}\to C_{n-1}$ defined by
$\partial_{n}[g_{1}|\ldots|g_{n}]=[g_{2}|\ldots|g_{n}]+\sum_{i=1}^{n-1}(-1)^{i}[g_{1}|\ldots|g_{i}g_{i+1}|\ldots|g_{n}]+(-1)^{n}[g_{1}|\ldots|g_{n-1}].$
Write $Z_{n}=\ker\partial_{n}$ and $B_{n}=\operatorname{im}\partial_{n+1}$. Then we define the $n$-th integral homology group of $G$ by $H_{n}(G,\mathbb{Z})=H_{n}(C_{\bullet})=Z_{n}/B_{n}$.
The differential $\partial_{n+1}$ induces the map $\delta^{n}:\operatorname{Hom}(C_{n},A)\to\operatorname{Hom}(C_{n+1},A)$. Denote by $Z^{n}=Z^{n}(G,A)=\ker\delta^{n}$ the set of $n$-cocycles, and by $B^{n}=B^{n}(G,A)=\operatorname{im}\delta^{n-1}$ the set of $n$-coboundaries. The $n$-th cohomology group of $G$ with coefficients in $A$ is defined by $H^{n}(G,A)=H^{n}(\operatorname{Hom}(C_{\bullet},A))=Z^{n}/B^{n}$.
Now we define a more general version of the $\operatorname{Utw}^{n}$ group by
$\operatorname{Utw}^{n}(G,A)=\{[\omega]\in H^{n}(G,A)\mid\sum_{\sigma\in S_{n}}\operatorname{sgn}(\sigma)\cdot\omega(g_{\sigma(1)},\dots,g_{\sigma(n)})=0\\ \hbox{ for all }(g_{1},\dots,g_{n})\in X_{n}(G)\}.$
Note that $\operatorname{Utw}^{n}(G)=\operatorname{Utw}^{n}(G,U(1))$. We prove the following version of Universal Coefficient Theorem [3, p. 60] for the $\operatorname{Utw}^{n}$ groups:
Theorem 4.1.
Let $G$ be a group and $A$ a trivial $G$-module. Then there is a short exact sequence
(4.1.1) $0\longrightarrow\operatorname{Ext}_{\mathbb{Z}}^{1}\!\bigl(H_{n-1}(G,\mathbb{Z}),A\bigr)\ \xrightarrow{\ \iota\ }\ \operatorname{Utw}^{n}(G,A)\ \xrightarrow{\ \tilde{\varrho}\ }\ \operatorname{Hom}\!\bigl(H_{0n}(G,\mathbb{Z}),A\bigr)\longrightarrow 0,$
with $H_{0n}(G,\mathbb{Z})$ being a quutient of $H_{n}(G,\mathbb{Z})$ defined as $H_{0n}(G,\mathbb{Z})=Z_{n}/(Z_{0n}+B_{n})$, where
$Z_{0n}=\left\langle\sum_{\sigma\in S_{n}}\operatorname{sgn}\sigma\cdot[g_{\sigma(1)}|\dots|g_{\sigma(n)}]\mid(g_{1},\dots,g_{n})\in X_{n}(G)\right\rangle.$
The sequence (4.1.1) splits (non-canonically).
Proof.
As $A$ is a trivial $G$-module, the Universal Coefficient Theorem (UCT) yields a short exact sequence [3, p. 60]
(4.1.2) $0\longrightarrow\operatorname{Ext}_{\mathbb{Z}}^{1}\!\bigl(H_{n-1}(G,\mathbb{Z}),A\bigr)\ \xrightarrow{\ \iota\ }\ H^{n}(G,A)\ \xrightarrow{\ \varrho\ }\ \operatorname{Hom}\!\bigl(H_{n}(G,\mathbb{Z}),A\bigr)\longrightarrow 0.$
that splits (non-canonically).
The maps $\iota$ and $\varrho$ in (4.1.2) can be described as follows. A cohomology class $\omega]\in H^{n}(G,A)$ can be represented with $\omega\in\operatorname{Hom}(C_{n},A)$ satisfying $\omega\partial_{n+1}=0$. Then we have a well-defined evaluation map $\varrho([\omega]):Z_{n}/B_{n}\to A$ given by $\varrho([\omega])(z+B_{n})=\omega(z)$. Specifically, if
$z=\sum_{k}\alpha_{k}[g_{1k}|\dots|g_{nk}],\,\hbox{ where }\alpha_{k}\in\mathbb{Z},$
then
$\varrho([\omega])(z+B_{n})=\sum_{k}\alpha_{k}\omega(g_{1k},\dots,g_{nk}).$
To define $\iota$, note that the long $\operatorname{Ext}$-sequence associated to $0\to B_{n-1}\to Z_{n-1}\to H_{n-1}\to 0$ gives an exact sequence
$\operatorname{Hom}(Z_{n-1},A)\xrightarrow{\ \operatorname{res}\ }\operatorname{Hom}(B_{n-1},A)\longrightarrow\operatorname{Ext}_{\mathbb{Z}}^{1}(H_{n-1},A)\longrightarrow 0.$
Thus, $\operatorname{Ext}_{\mathbb{Z}}^{1}(H_{n-1},A)$ can be identified with the quotient of $\operatorname{Hom}(B_{n-1},A)$ modulo the image of the restriction map $\operatorname{res}:\operatorname{Hom}(Z_{n-1},A)\to\operatorname{Hom}(B_{n-1},A)$. With this identification, the map $\iota$ is given by $\iota(\varphi+\operatorname{im}\operatorname{res})=[\varphi\partial_{n}]$.
At first note that $Z_{0n}$ is contained in $Z_{n}$ by [3, p. 36]. Take an arbitrary element of $\operatorname{Ext}_{\mathbb{Z}}^{1}\!\bigl(H_{n-1}(G,\mathbb{Z}),A\bigr)$ and represent it as $\varphi+\operatorname{im}\operatorname{res}$ as above. Then $\iota(\varphi+\operatorname{im}\operatorname{res})=[\varphi\partial_{n}]\in H^{n}(G,A)$. We claim that $\varphi\partial_{n}]$ actually belongs to $\operatorname{Utw}^{n}(G,A)$. To this end, take an arbitrary commuting $n$-tuple $(g_{1},\dots,g_{n})\in X_{n}(G)$. Then
$\sum_{\sigma\in S_{n}}\operatorname{sgn}\sigma\cdot\varphi\partial_{n}(g_{\sigma(1)},\dots,g_{\sigma(n)})=\varphi\partial_{n}\left(\sum_{\sigma\in S_{n}}\operatorname{sgn}\sigma\cdot[g_{\sigma(1)}|\dots|g_{\sigma(n)}]\right)=0,$
as required.
Given $\omega]\in\operatorname{Utw}^{n}(G,A)$ and a commuting tuple $(g_{1},\dots,g_{n})$, we note that
$\varrho([\omega])\left(\sum_{\sigma\in S_{n}}\operatorname{sgn}\sigma\cdot[g_{\sigma(1)}|\dots|g_{\sigma(n)}]+B_{n}\right)=\sum_{\sigma\in S_{n}}\operatorname{sgn}\sigma\cdot\omega(g_{\sigma(1)},\dots,g_{\sigma(n)})=0.$
This shows that $\varrho([\omega])$ induces a homomorphism $H_{0n}(G,\mathbb{Z})\to A$. It follows that $\varrho$ induces an epimorphism $\tilde{\varrho}:\operatorname{Utw}^{n}(G,A)\to\operatorname{Hom}(H_{0n}(G,\mathbb{Z}),A)$ such that the following diagram commutes:
${H^{n}(G,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varrho}$${\operatorname{Hom}(H_{n}(G,\mathbb{Z}),A)}$${\operatorname{Utw}^{n}(G,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{\varrho}}$${\operatorname{Hom}(H_{0n}(G,\mathbb{Z}),A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi^{*}}$
Here $\pi^{*}$ is induced by the epimorphism $\pi:H_{n}(G,\mathbb{Z})\to H_{0n}(G,\mathbb{Z})$. It is easy to see that $\ker\tilde{\varrho}=\operatorname{im}\iota$. This concludes the proof by the UCT. ∎
As $U(1)$ is a divisible abelian group, we have that $\operatorname{Ext}_{\mathbb{Z}}^{1}(H_{n-1}(G,\mathbb{Z}),U(1))=0$ for a finite group $G$. Note that the UCT, in particular, implies that if $G$ is a finite group, then $H^{n}(G,U(1))\cong\operatorname{Hom}(H_{n}(G,\mathbb{Z}),U(1))$. In the same way, Theorem 4.1 gives that $\operatorname{Utw}^{n}(G)\cong\operatorname{Hom}(H_{0n}(G,\mathbb{Z}),U(1)).$ This yields an isomorphism
$\operatorname{Utw}^{n}(G)\cong H_{0n}(G,\mathbb{Z}).$
This can be applied to compute the abelian invariants of $\operatorname{Utw}^{n}(G)$ using the functionality of HAP [10]. We elaborate on the algorithm in Appendix A.
Another way of computing $\operatorname{Utw}^{n}(G)$ would be to directly compute the untwisted cocycles. This is also described in Appendix A, and the code is available [23]. This also provides a way of computing the DW partition function explicitly. It should be noted, however, that HAP currently only supports 2-cocycles and 3-cocycles. We note that HAP already has a built-in function DijkgraafWittenInvariant that computes the DW partition function for a given $3$-manifold $M$ and a finite group $G$.
As in the case of 2D discrete torsion, it is evident that $G$ admits untwisted disrete torsion of dimension $n$ if and only if $H_{0n}(G,\mathbb{Z})$ is non-trivial. Also, $G$ is $n$-untwisted if and only if $H_{0n}(G,\mathbb{Z})=H_{n}(G,\mathbb{Z})\neq 0$.
Given a group $G$, define
$_{n}(G)=H_{n}(G,\mathbb{Z})/I_{n}(G),$
where $I_{n}(G)$ is the subgroup of $H_{n}(G,\mathbb{Z})$ generated by all the images of the induced maps $H_{n}(B,\mathbb{Z})\to H_{n}(G,\mathbb{Z})$, where $B$ runs through all abelian subgroups of $G$. By a standard duality argument, we have that $\operatorname{Hom}({}_{n}(G),U(1))$ is isomorphic to
$^{n}(G,U(1))=\ker\operatorname{res}:H^{n}(G,U(1))\to\bigoplus_{\begin{subarray}{c}B\leq G\\ B\hbox{\tiny abelian}\end{subarray}}H^{n}(B,U(1)),$
[…]
Theorem 4.2.
[…]
Proof.
[…]
$\prod_{\sigma\in S_{n}}\omega(g_{\sigma(1)},\dots,g_{\sigma(n)})^{\operatorname{sgn}\sigma}=1.$
As this holds for every $\mathbf{g}\in X_{n}(G)$, we deduce that $\omega]\in\operatorname{Utw}^{n}(G)$. ∎
[…]
5. Examples
In this section we report on some computational examples that were obtained. All the calculations were perfomed on Macbook M1 Pro with 16 GB RAM. The code is publicly available [23].
5.1. Untwisted groups
We have showed in Section 3 that finite subgroups of $\operatorname{SU}(3)$ have trivial Bogomolov multipliers. We can use the computational tools to deal with the finite subgroups of $SU(4)$ described in [11]. We compute the second integral homology and Bogomolov multipliers of the groups I* – XXX*. [^1] We point out that table of the groups XXII* – XXX* in [11] contains a typo; The matrices $R$, $S$ and $T$ should be replaced by $R^{\prime}$, $S^{\prime}$ and $T^{\prime}$, respectively. It turns out that these groups always have trivial $\operatorname{Utw}^{2}$.
We have also performed some calculations with groups XXXI* – XXXIV*, which are actually families depending on parameters. [^2] Note that [11] contains two families with the same name XXXIII*. For lots of choices of these, the Bogomolov multiplier is always trivial. We have not found any case with $\operatorname{Utw}^{2}(G)$ non-trivial for a finite subgroup $G$ of $\operatorname{SU}(4)$, and we conjecture this never happens.
We list some cases of untwisted groups, that is, groups $G$ with $\operatorname{Utw}^{2}(G)=H^{2}(G,U(1))\neq 1$. In GAP [9] notation for small groups, the smallest such group is $[64,182]$ (182182nd group in the library of all groups of order $64$). This is one of the eight groups of order 64 with non-trivial Bogomolov multiplier. There are two groups of order $3^{5}$ that are untwisted, namely $[243,29]$ and $[243,30]$. There are 5 untwisted groups of order $5^{5}$. These are $[3125,34]$, $[3125,35]$, $[3125,36]$, $[3125,37]$, $[3125,38]$. Similarly, the groups $[16807,38]$, $[16807,39]$, $[16807,40]$, $[16807,41]$ and $[16807,42]$ are the utwisted groups of order $7^{5}$. We collect some statistics on other untwisted groups in Table 2.
Order Nr all Nr untwisted
$2^{7}$ 2328 16
$2^{6}\cdot 3$ 1543 6
$2^{6}\cdot 5$ 1640 6
$2^{6}\cdot 7$ 1396 6
$2^{6}\cdot 3^{2}$ 8681 26
$3^{6}$ 504 14
$3^{6}\cdot 2$ 1798 46
$3^{7}$ 9310 153
$5^{6}$ 684 38
Table 2. Untwisted groups.
5.2. Higher dimensional discrete torsion
Basic examples
We first consider the cyclic group $\mathbb{Z}/m$ of order $m$. It is known that $H_{n}(\mathbb{Z}/m,\mathbb{Z})$ is zero for even $n>0$, and isomorphic to $\mathbb{Z}/m$ for all odd $n$. By definition, it is straighforward to verify that $H_{0n}(\mathbb{Z}/m,\mathbb{Z})=H_{n}(\mathbb{Z}/m,\mathbb{Z})$ for all $n>0$. This shows that cyclic groups are $n$-untwisted for all odd $n$.
Consider the group $(\mathbb{Z}/2)^{3}$. The calculations are given in the Table 3. We observe that $(\mathbb{Z}/2)^{3}$ is not $3$-untwisted, yet it is both $4$-untwisted and $5$-untwisted. The results also show, for example, that $\operatorname{Utw}^{3}$ does not in general commute with direct products, as opposed to $\operatorname{Utw}^{2}$ [16].
$n$ $H_{n}(G,\mathbb{Z})$ $H_{0n}(G,\mathbb{Z})$ $H_{n}$ time [ms] $H_{0n}$ time [ms]
2 $(\mathbb{Z}/2)^{3}$ 0 15 14
3 $(\mathbb{Z}/2)^{7}$ $(\mathbb{Z}/2)^{6}$ 13 121
4 $(\mathbb{Z}/2)^{8}$ $(\mathbb{Z}/2)^{8}$ 18 5966
5 $(\mathbb{Z}/2)^{13}$ $(\mathbb{Z}/2)^{13}$ 27 507232
Table 3. Computations for $G=(\mathbb{Z}/2)^{3}$.
Next we consider symmetric groups. These have trivial Bogomolov multiplier [19], so we focus on the dimensions 3 and 4, as the Bogomolov multipliers of alternating groups are always trivial [19]. Table 4 Table 5 indicate that small symmetric groups are not $3$-untwisted, but they are $4$-untwisted. The alternating groups behave quite differently, similar computations show that $A_{m}$, $3\leq m\leq 7$ are $3$-untwisted. On the other hand, $H_{4}(A_{m},\mathbb{Z})=0$ for $m\leq 7$; the group $A_{8}$ is the smallest alternating group with non-trivial $4$-th homology. Note that $\operatorname{Utw}^{2}(A_{m})$ is trivial by [19].
Computations with dihedral groups $D_{2m}$, $m\leq 12$, show that these groups are $3$-untwisted, and $4$-untwisted when $m$ is even; note that $H_{4}(D_{2m},\mathbb{Z})=0$ for $m$ odd. Bogomolov multipliers of dihedral groups are always trivial by Proposition 3.3. If $m\leq 6$, the computations show that $D_{2m}$ is $5$-untwisted. The calculations can be pushed to show that $D_{8}$ is $6$-untwisted (around 7 seconds of CPU time) and $7$-untwisted (30 minutes of CPU time).
$m$ $H_{3}(S_{m},\mathbb{Z})$ $H_{03}(S_{m},\mathbb{Z})$ $H_{3}$ time [ms] $H_{03}$ time [ms]
3 $\mathbb{Z}/6$ $\mathbb{Z}/6$ 1 75
4 $\mathbb{Z}/6\oplus\mathbb{Z}/4$ $\mathbb{Z}/6\oplus\mathbb{Z}/4$ 30 29
5 $\mathbb{Z}/6\oplus\mathbb{Z}/4$ $\mathbb{Z}/6\oplus\mathbb{Z}/4$ 42 400
6 $\mathbb{Z}/2\oplus\mathbb{Z}/4\oplus\mathbb{Z}/6$ $\mathbb{Z}/4\oplus\mathbb{Z}/6$ 84 6077
7 $\mathbb{Z}/2\oplus\mathbb{Z}/4\oplus\mathbb{Z}/6$ $\mathbb{Z}/4\oplus\mathbb{Z}/6$ 123 123878
Table 4. Computations for $G=S_{m}$, dimension 3.
$m$ $H_{4}(S_{m},\mathbb{Z})$ $H_{04}(S_{m},\mathbb{Z})$ $H_{4}$ time [ms] $H_{04}$ time [ms]
3 0 0 2 25
4 $\mathbb{Z}/2$ $\mathbb{Z}/2$ 21 361
5 $\mathbb{Z}/2$ $\mathbb{Z}/2$ 24 6187
6 $(\mathbb{Z}/2)^{2}$ $(\mathbb{Z}/2)^{2}$ 95 103905
Table 5. Computations for $G=S_{m}$, dimension 44.
Finite subgroups of $SU(4)$
We now focus on some finite subgroups $G$ of $SU(4)$ which are listed in [11]. For these, we compute $H_{3}(G,\mathbb{Z})$ and $H_{03}(G,\mathbb{Z})$. We collect the results in Table 6. The table is not complete due to performance obstacles, see Appendix A.
The computations show a startling difference between dimensions 2 and 3. While there are no known subgroups $G$ of $SU(4)$ with non-trivial Bogomolov multiplier, there several cases that are $3$-untwisted. The group $\operatorname{Utw}^{3}(G)$ is usually non-trivial and very close to $H^{3}(G,U(1))$.
$G$ $|G|$ $H_{3}(G,\mathbb{Z})$ $H_{03}(G,\mathbb{Z})$
I* 240 $\mathbb{Z}/30\oplus\mathbb{Z}/8$ $\mathbb{Z}/30\oplus\mathbb{Z}/8$
II* 60 $\mathbb{Z}/30$ $\mathbb{Z}/30$
VIII* 480 $\mathbb{Z}/12\oplus\mathbb{Z}/8$ $\mathbb{Z}/12\oplus\mathbb{Z}/8$
X* 288 $(\mathbb{Z}/3)^{2}\oplus\mathbb{Z}/6\oplus\mathbb{Z}/8$ $(\mathbb{Z}/3)^{2}\oplus\mathbb{Z}/6\oplus\mathbb{Z}/4$
XXXI*(1) 12 $\mathbb{Z}/6$ $\mathbb{Z}/6$
XXXI*(2) 96 $(\mathbb{Z}/2)^{3}\oplus\mathbb{Z}/3\oplus\mathbb{Z}/8$ $(\mathbb{Z}/2)^{3}\oplus\mathbb{Z}/3\oplus\mathbb{Z}/4$
XXXI*(3) 324 $\mathbb{Z}/2\oplus(\mathbb{Z}/3)^{3}$ $\mathbb{Z}/2\oplus(\mathbb{Z}/3)^{2}$
XXXII*(1) 24 $\mathbb{Z}/2\oplus\mathbb{Z}/3\oplus\mathbb{Z}/4$ $\mathbb{Z}/2\oplus\mathbb{Z}/3\oplus\mathbb{Z}/4$
XXXII*(2) 192 $(\mathbb{Z}/2)^{2}\oplus\mathbb{Z}/3\oplus(\mathbb{Z}/4)^{3}$ $(\mathbb{Z}/2)^{2}\oplus\mathbb{Z}/3\oplus(\mathbb{Z}/4)^{3}$
XXXIII*(1) 8 $(\mathbb{Z}/2)^{2}\oplus\mathbb{Z}/4$ $(\mathbb{Z}/2)^{2}\oplus\mathbb{Z}/4$
XXXIII*(2) 64 $(\mathbb{Z}/2)^{5}\oplus(\mathbb{Z}/4)^{3}$ $(\mathbb{Z}/2)^{4}\oplus(\mathbb{Z}/4)^{3}$
XXXIII*(3) 216 $(\mathbb{Z}/6)^{2}\oplus\mathbb{Z}/4$ $(\mathbb{Z}/6)^{2}\oplus\mathbb{Z}/4$
XXXIII*(1) 4 $(\mathbb{Z}/2)^{3}$ $(\mathbb{Z}/2)^{3}$
XXXIII*(2) 32 $(\mathbb{Z}/2)^{9}\oplus\mathbb{Z}/8$ $(\mathbb{Z}/2)^{9}\oplus\mathbb{Z}/4$
XXXIII*(3) 108 $(\mathbb{Z}/2)^{3}\oplus(\mathbb{Z}/3)^{4}$ $(\mathbb{Z}/2)^{3}\oplus(\mathbb{Z}/3)^{3}$
Table 6. Computations for $G\leq\operatorname{SU}(4)$, dimension 33.
[…]
6. Conclusions
We introduced and studied the groups $\operatorname{Utw}^{n}(G)\subseteq H^{n}(G,U(1))$ consisting of cohomology classes whose Dijkgraaf–Witten weights are trivial on all commuting nn–tuples, i.e. on all flat GG–bundles over the nn–torus. Conceptually, $\operatorname{Utw}^{n}(G)$ captures the part of group cohomology that is invisible to torus backgrounds. Operationally, it measures the failure of $H^{n}(G,U(1))$ to parametrize physically distinct discrete-torsion deformations at the level of torus amplitudes. For finite $G$ we identified $\operatorname{Utw}^{n}(G)$ with $\operatorname{Hom}(H_{0n}(G,\mathbb{Z}),U(1))$ and hence with an explicit quotient $H_{0n}(G,\mathbb{Z})$ of $H_{n}(G,\mathbb{Z})$, providing a purely homological handle on this “torus-invisible” sector.
In degree $2$ this recovers the familiar Bogomolov multiplier/unramified Brauer group and its efficient dual description. Our computations indicate a rigidity for the finite subgroups of $\operatorname{SU}(4)$ in the Hanany–He classification [11]. Despite non-trivial Schur multipliers occurring frequently, the Bogomolov multipliers we tested are trivial, so there is no genuine two-dimensional discrete torsion visible on the torus in these examples. In higher degrees the behavior is richer. The groups $H_{0n}(G,\mathbb{Z})$ need not coincide with $H_{n}(G,\mathbb{Z})$, and we observe examples where untwistedness depends sensitively on $n$. This highlights that higher-dimensional discrete torsion is not merely a repackaging of ordinary group cohomology, but is constrained by the geometry of commuting tuples.
[…]
Appendix A Algorithms
A.1. Algorithm for conjugacy class representatives of commuting tuples
Let $G$ be a finite group. Recall that the set of all commuting $n$-tuples is
$X_{n}(G)=\{(g_{1},\dots,g_{n})\in G^{n}\mid g_{i}g_{j}=g_{j}g_{i}\text{ for all }i,j\}.$
The group $G$ acts on $X_{n}(G)$ by diagonal conjugation,
$k\cdot(g_{1},\dots,g_{n})=(kg_{1}k^{-1},\dots,kg_{n}k^{-1}),\qquad k\in G.$
We seek a set of orbit representatives for this action, i.e. a transversal of $X_{n}(G)/G$, the moduli space of flat connections over $T^{n}$. The representatives are obtained recursively using a centralizer-based search, which prunes the space of possible tuples efficiently. We choose an element $g_{1}\in G$ and restrict to its centralizer $C_{G}(g_{1})$. Then we choose $g_{2}\in C_{G}(g_{1})$ and restrict to $C_{G}(g_{1},g_{2})$. Continue this process until an $n$-tuple $\mathbf{g}=(g_{1},\dots,g_{n})$ is obtained. Each constructed tuple automatically satisfies $g_{i}g_{j}=g_{j}g_{i}$ for all $i,j$. Then we apply the orbit–stabilizer algorithm to eliminate duplicates under the conjugation action of $G$, keeping one representative per orbit. For each orbit representative $\mathbf{g}$, its stabilizer subgroup
$C_{G}(\mathbf{g})=\{k\in G\mid kg_{i}k^{-1}=g_{i}\ \forall i\}$
is computed. Its size $|C_{G}(\mathbf{g})|$ follows from the orbit–stabilizer theorem and serves as a natural weight in topological applications.
At each step the search space is reduced to a smaller centralizer, so the algorithm avoids testing all $|G|^{n}$ possible tuples. In particular, the search depth is $n$, and at each level only elements commuting with all previously chosen ones are explored. The final reduction by conjugation further eliminates equivalent configurations.
A.2. The dual of the untwisted group
We now describe an algorithm for computing the isomorphism type of the group $H_{0n}(G,\mathbb{Z})$.
Let $C_{\bullet}$ be a chain complex of free abelian groups
$\cdots\xrightarrow{\partial_{n+1}}C_{n}\xrightarrow{\partial_{n}}C_{n-1}\xrightarrow{}\cdots,\qquad C_{k}\cong\mathbb{Z}^{r_{k}},$
and let $Z_{0n}\leq C_{n}$ be a subgroup generated by a finite list of vectors $v^{(1)},\dots,v^{(s)}\in\mathbb{Z}^{r_{n}}$; these are the images of alternating bar words under a chain map. Recall that $B_{n}=\mathrm{im}(\partial_{n+1})\leq C_{n}$ and $Z_{n}:=\ker(\partial_{n})\leq C_{n}$. We need to compute the quotient abelian group
$Q_{n}\;:=\;Z_{n}/(B_{n}+Z_{0n}),$
where
$Z_{0n}=\left\langle\sum_{\sigma\in S_{n}}\operatorname{sgn}\sigma\cdot[g_{\sigma(1)}|\dots|g_{\sigma(n)}]\mid(g_{1},\dots,g_{n})\in X_{n}(G)\right\rangle.$
Step 1: boundary matrices. Represent $\partial_{n}$ and $\partial_{n+1}$ by integer matrices
$D_{n}\in M_{r_{n-1}\times r_{n}}(\mathbb{Z}),\qquad D_{n+1}\in M_{r_{n}\times r_{n+1}}(\mathbb{Z}),$
with respect to the standard bases of $C_{k}\cong\mathbb{Z}^{r_{k}}$.
Step 2: a $\mathbb{Z}$-basis of $Z_{n}=\ker(D_{n})$. Compute the Smith normal form of $D_{n}$, i.e. unimodular matrices $U\in\mathrm{GL}_{r_{n-1}}(\mathbb{Z})$ and $V\in\mathrm{GL}_{r_{n}}(\mathbb{Z})$ such that
$UD_{n}V\;=\;S\;=\;\mathrm{diag}(s_{1},\dots,s_{r},0,\dots,0),\qquad r=\mathrm{rank}(D_{n}).$
Then $\ker(S)=\{0\}^{r}\times\mathbb{Z}^{r_{n}-r}$, hence the last $k:=r_{n}-r$ columns of $V$ form a $\mathbb{Z}$-basis of $Z_{n}=\ker(D_{n})$.
Equivalently, for any $z\in Z_{n}$ the coordinate vector $y:=V^{-1}z$ satisfies $y_{1}=\cdots=y_{r}=0$, and the tail
$\mathrm{coord}(z)\;:=\;(y_{r+1},\dots,y_{r_{n}})\in\mathbb{Z}^{k}$
records the coordinates of $z$ in the chosen kernel basis.
Step 3: relations coming from $B_{n}+Z_{0n}$. Let $L:=B_{n}+Z_{0n}\leq Z_{n}$. A generating set for $L$ is obtained by taking: (i) the columns of $D_{n+1}$ (generators of $B_{n}$), and (ii) the prescribed generators $v^{(j)}$ of $Z_{0n}$. For each generator $\ell\in L$, compute its kernel coordinates
$c(\ell)\;:=\;\mathrm{coord}(\ell)\in\mathbb{Z}^{k}.$
Collect these coordinate vectors as columns of an integer matrix
$A\in M_{k\times m}(\mathbb{Z}),$
so that $\mathrm{im}(A)\subseteq\mathbb{Z}^{k}$ is the subgroup of relations in kernel coordinates.
Step 4: quotient invariants. By construction,
$Q_{n}\;\cong\;\mathbb{Z}^{k}\big/\mathrm{im}(A).$
Compute the Smith normal form of $A$:
$U^{\prime}AV^{\prime}\;=\;\mathrm{diag}(d_{1},\dots,d_{t},0,\dots,0),\quad d_{i}\geq 2,\quad d_{i}\mid d_{i+1}.$
It follows that
$Q_{n}\;\cong\;\bigoplus_{i=1}^{t}\mathbb{Z}/d_{i}\mathbb{Z}\ \oplus\ \mathbb{Z}^{\,k-t}.$
In particular, the elementary divisors $d_{i}$ (together with the free rank ktk-t) give the complete isomorphism type of $Z_{n}/(B_{n}+Z_{0n})$ as a finitely generated abelian group.
The advantage of this algorithm is that it employs integer linear algebra to avoid costly coset enumeration used in computing the quotients. The computational bottleneck is the construction and use of the chain equivalence $\phi_{n}:BC_{n}(G)\to T_{n}=\mathbb{Z}\otimes_{\mathbb{Z}G}R_{n}$, where $R$ is a free $\mathbb{Z}G$-resolution produced by HAP. The chain map $\phi_{*}$ can be obtained via HAP’s commands BarComplexEquivalence or BarResolutionEquivalence. Applying $\phi_{n}$ to many alternating bar words forces repeated, expensive reductions through the chosen $R$, and this dominates runtime. By contrast, generating commuting tuple representatives and the final Smith-normal-form step for $Z_{n}/(B_{n}+Z_{0n})$ are typically negligible here. The timings indicate that the cost is overwhelmingly in evaluating $\phi_{n}$ (and in practice, largely in building the underlying equivalence data and rewriting words in $R_{n}$).
Regarding the generating set of $\mathbb{Z}_{0n}$, note that $C_{\bullet}$ is the complex of coinvariants $(F_{\bullet})_{G}$, where $F_{\bullet}$ is the bar resolution of $\mathbb{Z}$. That implies $1\otimes x=1\otimes(g\cdot x)$ for all $g\in G$ and $x\in F_{\bullet}$. The action of $G$ on $F_{\bullet}$ is conjugation, therefore, in our notation, $hg_{1}h^{-1}|\dots|hg_{n}h^{-1}]=[g_{1}|\dots|g_{n}]$ for all $h,g_{1},\dots,g_{n}\in G$. Therefore,
$Z_{0n}=\left\langle\sum_{\sigma\in S_{n}}\operatorname{sgn}\sigma\cdot[g_{\sigma(1)}|\dots|g_{\sigma(n)}]\mid(g_{1},\dots,g_{n})\in X_{n}(G)/G\right\rangle.$
This, in general, yields a much smaller set of generators than running through all commuting $n$-tuples in $G$.
A.3. Computation of cocycles with values in $U(1)$
In order to deal with the explicit elements in $H^{n}(G,U(1))$ and compute the phase factors, one needs to find a set of representatives of $n$-cocoycles that define $\operatorname{Utw}^{n}(G)$. These are maps $G^{n}\to U(1)$. The first obstacle is that GAP does not provide functionality for computing with complex numbers in general. It does, however, handle arithmetics with roots of unity. It turns out that this suffices for our purposes. Namely, If $G$ is a finite group, Then $H^{n}(G,U(1))$ is canonically isomorphic to $\operatorname{Hom}(H_{n}(G,\mathbb{Z}),U(1))$. Noter that $H_{n}(G,\mathbb{Z})$ is a finite abelian group. Let $m$ be a positive number with $m\cdot H_{n}(G,\mathbb{Z})=0$. Then the above isomorphism implies that every $n$-cocycle $\omega\in Z^{n}(G,U(1))$ actually maps into the subgroup of $U(1)$ generated by the primitive $m$-th root of unity $\zeta_{m}=e^{2\pi i/m}$. Thus, one can replace $U(1)$ with the cyclic group $\mathbb{Z}/m$ of order $m$ to perform the computations with the cocycles.
While HAP computes $H^{n}(G,A)$ for finite abelian coefficient modules $A$, access to cochain-level data (explicit cocycles and their manipulation) is currently provided only for degrees $n\leq 3$. Thus, we are currently limited to describing the elements of $\operatorname{Utw}^{2}(G)$ and $\operatorname{Utw}^{3}(G)$. The algorithm and its implementation that we present are nevertheless general enough to cover higher degrees once the corresponding cocycle-level functionality becomes available.
Once we have computed the cocycles with values in $\mathbb{Z}/m$, one can obtain their classes in $H^{n}(G,\mathbb{Z}/m)$. This group may not be isomorphic to $H^{n}(G,U(1))$. To make up for that, note that the UCT implies that
$H^{n}(G,\mathbb{Z}/m)\cong\operatorname{Ext}_{\mathbb{Z}}^{1}(H_{n-1}(G,\mathbb{Z}),\mathbb{Z}/m)\oplus\operatorname{Hom}(H_{n}(G,\mathbb{Z}),\mathbb{Z}/m).$
The second summand is isomorphic to $H_{n}(G,\mathbb{Z})$, as $m\cdot H_{n}(G,\mathbb{Z})=0$. Therefore,
$H^{n}(G,\mathbb{Z}/m)\cong\operatorname{Ext}_{\mathbb{Z}}^{1}(H_{n-1}(G,\mathbb{Z}),\mathbb{Z}/m)\oplus H^{n}(G,U(1)).$
A similar argument shows that Theorem 4.1 implies
$\operatorname{Utw}^{n}(G,\mathbb{Z}/m)\cong\operatorname{Ext}_{\mathbb{Z}}^{1}(H_{n-1}(G,\mathbb{Z}),\mathbb{Z}/m)\oplus\operatorname{Utw}^{n}(G).$
These show that $H^{n}(G,U(1))$ and $\operatorname{Utw}^{n}(G)$ can be obtained by subtracting the direct summand $\operatorname{Ext}_{\mathbb{Z}}^{1}(H_{n-1}(G,\mathbb{Z}),\mathbb{Z}/m)$ from $H^{n}(G,\mathbb{Z}/m)$ and $\operatorname{Utw}^{n}(G,\mathbb{Z}/m)$. The isomorphism type of the $\operatorname{Ext}$ part follows from the formula
$\operatorname{Ext}_{\mathbb{Z}}^{1}\left(\bigoplus_{i}\mathbb{Z}/d_{i},\mathbb{Z}/m\right)\cong\bigoplus_{i}\mathbb{Z}/\gcd(d_{i},m).$
A.4. Connection with Dijkgraaf–Witten theory
In the Dijkgraaf–Witten topological gauge theory for a finite group $G$ and cocycle $\omega\in H^{n}(G,U(1))$, field configurations on a closed $n$-manifold $M$ correspond to flat $G$-bundles, classified by group homomorphisms
$\phi\colon\pi_{1}(M)\to G,$
up to conjugation. For the $n$-torus $T^{n}$, one has $\pi_{1}(T^{n})\cong\mathbb{Z}^{n}$, and therefore $\mathrm{Hom}(\pi_{1}(T^{n}),G)$ can be identified with $X_{n}(G)$, the set of commuting $n$-tuples in $G$. Two tuples related by conjugation represent gauge-equivalent bundles, so the physically distinct configurations are precisely the elements of the orbit space $X_{n}(G)/G$.
The Dijkgraaf–Witten partition function on $T^{n}$ takes the form
$Z^{\mathrm{DW}}_{\omega}(T^{n})\;=\;\sum_{[\mathbf{g}]\in X_{n}(G)/G}\frac{W_{\omega}(\mathbf{g})}{|C_{G}(\mathbf{g})|},$
and there is a similar expression for $Z^{\mathrm{orb}}_{\omega}(T^{n})$. Here $W_{\omega}(\mathbf{g})$ again denotes the phase obtained by evaluating on the tuple $\mathbf{g}$. The recursive centralizer algorithm described above provides an explicit finite enumeration of all orbit representatives $\mathbf{g}$ and stabilizer sizes $|C_{G}(\mathbf{g})|$, which are precisely the ingredients required to evaluate this partition function.